The temperature at which the speed of the sound in air becomes double of its value at 27 degree celsius is

Grade:

1 Answers

Arun
5 years ago
Use the formula 
velocity of sound waves v = sqrt (P/rho) 

P is the pressure and rho is the density.Write this as 

v1 = root(p1/rho) 
v2= root (p2/rho) 

square both sides 

v1^2 = p1 /rho 

v2^2 = p2/rho 

Take the ration 

v1^2/v2^2 = p1/p2 

Put the values v1=1 and v2 = 2 

YOu get 

1/4 = p1/p2 

p1 = n kb T1 and p2= n kb T2. Here n is the number density. kb Boltzmann constant and T1 and T2 are temperatures. 

If you use it in the above equation you get 

T2= 4 T1 

T1 is 27 degrees= 273 + 27 = 300 

T2= 1200 degrees = 1200-273 = 927 degrees.